OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)?

OpenStudy (photon336):

well the point slope form is like this \[y-y_1 = m(x-x_{1})\]

OpenStudy (anonymous):

A)y=2(x+2) B)y=2(x-2) C)y=-2(x-2) D)y=-2(x+2)

OpenStudy (anonymous):

i know, that's why im confused. Because the answers don't look like that. @Photon336

OpenStudy (photon336):

well first we can find the slope between the two points (-2,0) and (2,8) \[\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} } = m \]

OpenStudy (anonymous):

m=2, right ?

OpenStudy (photon336):

\[\frac{0-8 }{-2-2 } = \frac{ -8 }{ -4 } = 2 \]

OpenStudy (anonymous):

yay! i did that part right (:

OpenStudy (photon336):

now we know that m the slope is 2 now let's pick a point let's say (-2,0)

OpenStudy (anonymous):

what do we do with it ?

OpenStudy (photon336):

We put it into this equation here: \[y-y_1 = m(x-x_{1})\]

OpenStudy (photon336):

m = 2, so now we have We put it into this equation here: \[y-y_1 = 2(x-x_{1})\]

OpenStudy (photon336):

can you plug in (-2,0) in the appropriate places?

OpenStudy (photon336):

@unicornbeans

OpenStudy (anonymous):

y-0=2(x+2) ?

OpenStudy (anonymous):

OOOOOOHHHHHH !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (photon336):

so that becomes y = 2(x+2) y = 2x+4 now let's test this. to see if it passes through our points. let's plug in (-2) 2(-2)+4 = 0 (-2,0) y = 2x+4 y =2(2)+4 = 8 (2,8) you can clearly see that this equation passes through both points.

OpenStudy (anonymous):

So, is y=2(x+2) the answer then ?

OpenStudy (photon336):

what's your justification? why would it be the answer based on everything we said?

rebeccaxhawaii (rebeccaxhawaii):